OpenStudy (anonymous):

\(\Huge\cal\color{blue}{Hello!}\) Find the side length of the cube with the given volume. A. 625 m B. 125 m C. 25 m D. 5 m

OpenStudy (godlovesme):

\(\huge\color{violet}{WELCOME~TO~OS~:)}\) :D @i_just_made_new_jan

OpenStudy (anonymous):

can u helo?

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

one sec i'll try :)

OpenStudy (anonymous):

thx

OpenStudy (godlovesme):

yw :3

OpenStudy (godlovesme):

v= a^3 125= a^3 if u take the cube root of 125 u will get 5 :)

OpenStudy (godlovesme):

to check ur answer u can do 5^3 and u will get 125 :)

OpenStudy (anonymous):

thx...

OpenStudy (godlovesme):

anytime :D